# Functional Near-Infrared Spectroscopy (fNIRS) Analysis of Interaction Techniques in Touchscreen-Based Educational Gaming

### v5

@@ Abstract.
Interaction system, touch-based interaction, stylus, functional Near-Infrared Spectroscopy (fNIRS), brain signal, hemodynamic response, touchscreens, educational games
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: 27th International Conference on Multimodal Interaction (ICMI 2025) ; October 13–17, 2025; Canberra, Australiaisbn: 978-1-4503-XXXX-X/2018/06ccs: Human-centered computing Human computer interaction (HCI)ccs: Applied computing Interactive learning environmentsccs: Applied computing Learning management systemsccs: Applied computing Computer-assisted instruction
Figure 1. The overview of the study: study setup, game environment, and findings. (a) Experimental setup: A participant is playing an educational game on a touchscreen laptop, with a functional near-infrared spectroscopy (fNIRS) device affixed to the forehead. This setup captures hemodynamic responses in the prefrontal cortex (PFC), with data acquisition managed through cognitive optical brain imaging studio software on a secondary laptop. (b) Educational Game Environment: Sample quiz questions with options. (c) Hemodynamic Response: A heat map shows that using the hand required significantly less oxygen flow during the task than using a stylus. (d) Cognitive effort: Hand requires significantly less cognitive effort to get almost the same behavioral performance compared to the stylus, which suggests that input modality affects the neural activities.
@@ 1. Introduction
Touch-based interaction is becoming a more powerful modality due to its resemblance to the traditional pen-and-paper experience.  (Elttayef, 2016; Li et al., 2013; Taranta et al., 2016). Incorporating touchscreens into educational games has changed the way we interact while playing. This offers easy touch and movement on the screen, and makes applications more user friendly  (Crescenzi-Lanna, 2020).
Traditional human-computer interaction evaluations, such as questionnaires and think-aloud protocols or user behavior metrics such as reaction times, have limitations, including susceptibility to ambiguity, social pressures, and disruption of the interaction  (Frey et al., 2016; Kosch et al., 2023; Pettersson et al., 2018; Yu et al., 2023). The real-time behavioral metrics such as an functional near-infrared spectroscopy (fNIRS) data offer quantitative insights  (Lloyd-Fox et al., 2010; Yücel et al., 2021) where the traditional methods fail to accurately represent users’ mental states as they are unable to discern between factors such as low concentration or task complexity  (Frey et al., 2016; Fairclough, 2009; Frey et al., 2013; Pike et al., 2012) and attention as well as motivation  (Pike et al., 2012).
fNIRS, a noninvasive optical imaging technology, measures the changes in concentrations of oxygenated hemoglobin ($\Delta HbO$) and deoxygenated hemoglobin in the prefrontal cortex (PFC) to evaluate hemodynamic responses  (Lloyd-Fox et al., 2010; Herff et al., 2014; Yücel et al., 2021; Shewokis et al., 2015; Getchell and Shewokis, 2023; Yu et al., 2023). The fNIRS information provides human computer interaction researchers with a powerful tool to explore the complex relationship between cognitive processes and interactive technologies  (Frey et al., 2016; Kosch et al., 2023; Pettersson et al., 2018). This method assumes that brain or neural activity leads to an increase in oxygenated blood flow, as indicated by increased $\Delta HbO$. This suggests that during task performance, it requires more cognitive effort  (Sharmin et al., 2024; Lloyd-Fox et al., 2010; Herff et al., 2014; Shewokis et al., 2015; Yücel et al., 2021; Kosch et al., 2023; Pettersson et al., 2018; Yu et al., 2023). This connection helps us to understand the brain mechanisms behind cognitive functions, and build technologies and therapies to optimize or monitor them  (Lloyd-Fox et al., 2010; Yücel et al., 2021). The oxygen flow to the PFC can also be used to measure relative neural efficiency, which quantifies cognitive effort and behavioral performance. It can also measure relative neural involvement, which illustrates the relationship among motivation, engagement, and performance  (Paas et al., 2003; Shewokis et al., 2015; Getchell and Shewokis, 2023).
@@ 3.1. Participants
Outliers were identified and removed based on the standard interquartile range (IQR) method ($Q1-1.5\times IQR$ or $Q3+1.5\times IQR$)(Dodge, 2008). If the average $\Delta HbO$ values of a participant fell outside the range, they were considered outliers. The final sample comprised 16 graduate students (11 female, 5 male), aged 23–32 years ($M=27.13\pm 2.5$). Thus, we removed two more outliers after data collection.
Participants were graduate students aged between 23 and 32 years (11F,$M=27.13\pm 2.5$). The participants had strong experience using their hands on touchscreens, but only a few were confident using a stylus ($18.75\%$). More than half had never used games as a learning tool, while others used them occasionally. In terms of graph understanding, only a small number (6.25%) felt somewhat confident.
@@ 3.2. Apparatus
In this study, we used two computers and an fNIRS device (Figure 1 (a)). We used an 18-channel fNIRS imager 2000S(fNIRS Device LLC, Potomac, MD, USA), four light emitters with wavelengths ($\lambda$) of 730-850 nm and an average power of $<1$ mW, sampled at 2 Hz. The detectors were separated by 2.5 cm, which resulted in a penetration depth of approximately 1.2 cm. The fNIRS sensor band was aligned to match the horizontal and vertical axes of the head with those of the band. Specifically, the sensor’s vertical axis was placed at the Fp1 and Fp2 locations delineated in the international 10 - 20 system of cerebral electrode placement. Data was converted to concentration changes using the modified Beer-Lambert law (mBLL). The educational quiz games were played on a Dell XPS touchscreen laptop (L1) [Intel (R) Core(TM) i7-7700HQ processor at 2.80 GHz]. This laptop supports both hand touch and stylus input with ten touch points. We used a desktop computer (L2) [Intel(R) Core(TM) i7-10700T processor at 2.00 GHz] connected to the fNIRS device. Cognitive optical brain imaging studio Software was installed on L2 to gather fNIRS data. fNIRSoft Software (Version 4.9) evaluated the data. A Python code was executed on L2 and synchronized with each task to signal the fNIRS device to differentiate between the quiz and rest periods.
@@ 4.1. Quantitative Analysis
    These values were converted to Z-scores to standardize the data. Our aim was better performance with less cognitive effort. We used the mean task performance score to calculate the standardized performance index ($P_{z}$) as shown in Equation 2. We also computed the inverse of the mean $\Delta HbO$ values to reflect reduced effort, and used this in calculating the standardized average oxygenated hemoglobin index ($M_{z}$), as shown in Equation 3 (AKS, 2025; Paas et al., 2003; Shewokis et al., 2015; Getchell and Shewokis, 2023; Koiler et al., 2022). In the X-axis, we plot the standardized average $\Delta HbO$ and on the Y-axis, we plot the standardized performance level $P_{z}$. Then, we used a Euclidean transformation from vector geometry. The equation for RNE in Equation 4 and RNI in Equation 5 represent projections of the vector onto two rotated coordinate axes: one axis captures efficiency (RNE line, running bottom-left to top-right) and the other captures involvement (RNI line, bottom-right to top-left). These rotated axes help to separate two cognitive dimensions: how effectively someone learns, and how deeply involved they are.
@@ 4.2. Self-reported Measures & Instruments
We used the System Usability Scale (SUS) (Brooke et al., 1996), ten questions with a 5-point Likert-scale ranging from ‘strongly disagree’ to ‘strongly agree.’ On a scale of 0 to 100, the final SUS score was determined (0–50 = not acceptable, 51–67 = poor, 68 = okay, 69–80 = good, 81–100 = exceptional) (Brooke et al., 1996). We also used the NASA-TLX to assess the mental workload of participants during learning procedures. This assessment has six questions to produce an overall workload rating of 0-100 (Hart and Staveland, 1988).
@@ 5.1.2. Hemodynamic Response
Figure 5. Heatmap of $\Delta HbO$ across four regions of the PFC: (a) hand condition and (b) stylus condition. The stylus exhibiting elevated $\Delta HbO$ signifies greater neural activity compared to the hand condition.
@@ 5.1.3. Relative Neural Efficiency (RNE) and Relative Neural Involvement (RNI)
Figure 6. Relative neural efficiency and involvement: (a) relative neural efficiency (RNE), where the hand condition has higher efficiency with lower cognitive effort, and (b) lower relative neural involvement (RNI) than using the stylus condition. The red dot represents the hand, and the green diamond represents the stylus condition. The cognitive effort is higher in the stylus although the performance level is almost the same.
@@ 6. Discussion
Our work investigated the impacts of input modalities between two conditions (hand and stylus) on neural activity, cognitive effort, and user experience in a controlled study. We discuss our research questions based on our findings and how these modalities affect user interaction and perception in educational games. The behavioral performance results suggest that input modality (hand vs. stylus) did not significantly affect learning outcomes in the touchscreen-based quiz task. This indicates that task performance and efficiency were comparable across modalities, which supports the idea that both hand and stylus are equally effective for simple tap-based educational interactions.
@@
RQ1 focused on the hemodynamic response in educational gaming with two input methods (hand and stylus). Higher $\Delta HbO$ suggests higher oxygen flow to the brain, which is the result of higher neural activity (Kerr et al., 2022; Sharmin et al., 2024). In our study, the change in blood oxygenation levels ($\Delta HbO$) showed a significant decrement with hand usage compared to the stylus. The higher $\Delta HbO$ while using a stylus suggests that handling a stylus can require more precise motor control, fine motor skills, and coordination than using the hand directly. Thus, it is observed that the type of input method may change the cognitive effort in educational game studies. According to the demographic data, most participants were not regular stylus users. While this unfamiliarity could have potentially influenced their performance by increasing cognitive effort, our data did not reflect any significant disruption. Both response time and performance scores did not show any significant differences. This indicates that prior exposure to the stylus did not negatively affect their task performance, particularly for simple actions like tapping. However, this also highlights a critical design consideration: learning materials and interactive tools should be designed in such a way that users invest their effort primarily in the learning process, not in struggling to operate the interface itself.
@@
RQ2 focused on observing cognitive effort by measuring relative neural efficiency (RNE) and relative neural involvement (RNI). The results suggest that using the hand required less cognitive effort, as reflected in higher RNE and lower RNI values. On the other hand, the stylus condition showed higher neural involvement, but the performance score was almost the same as the hand condition. In the stylus condition, participants had to manage both holding the stylus and responding to quiz questions simultaneously, which required more tasks and led to more cognitive effort (Schwarz et al., 2014; Koiler et al., 2022; Getchell and Shewokis, 2023). This suggests that hands are more natural or intuitive in educational game play and also highlights the importance of interaction methods that support efficient performance with lower cognitive effort.
@@
RQ3 focused on the impact of self-reported usability and task load measures of different input methods in playing educational games. The SUS score showed that users found the hand-based input more intuitive and usable, with a good usability score. The NASA-TLX results indicated that the hand condition required less effort and caused less frustration. Together, these findings suggest that hand-based input is more efficient and less demanding cognitively than stylus-based interaction for educational gameplay.
@@
According to the results of RQ1, RQ2, and RQ3, the hand condition provides better outcomes than the stylus in both objective and subjective evaluations. Lower $\Delta HbO$ levels with hand usage are revealed by hemodynamic data (RQ1), which suggests less cognitive load. This is further supported by cognitive effort measurements (RQ2). These results suggest that the natural and intuitive usage of the hand leads to higher efficiency (RNE) and lower involvement (RNI) than stylus usage. These findings are supported by self-reported measures (RQ3), as NASA-TLX results show less task load and annoyance, and SUS ratings show improved usability for hand input. According to these findings, it might be easier to use the hand as an input method in educational games. The degree to which subjective impressions and objective facts agree demonstrates the importance of taking both viewpoints into account when assessing interaction techniques. This highlights the possibilities of hand-based input in learning games and provides information for developing systems that try to reduce cognitive effort while increasing effectiveness and user satisfaction.
@@
Our findings offer broader implications for the design of educational technology and natural user interfaces. The hand condition’s reduced cognitive effort indicates that familiarity and intuitiveness are critical for user experience and performance. These results contribute to human-computer interaction research, emphasizing that neuronal efficiency can be enhanced by reduced input complexity, particularly in educational settings. Additionally, this study demonstrates that neurophysiological measurements, like fNIRS, can provide objective validation for design choices that are often made primarily on subjective input. By highlighting the cognitive load differences between stylus and hand inputs, our work supports a user-centered approach to input modality selection. This neuro-informed perspective can be examined in future studies involving different input methods, such as mouse and interface designs. It may help develop adaptive educational tools that align interaction styles with users’ cognitive and behavioral performance.
@@
Limitations and Future Work: This study may have been limited by its small sample size. Differences in participants’ age, education, and background could have influenced the results. In future studies, we plan to recruit a more diverse and balanced participant group. A tablet or iPad could provide a more intuitive setup. Future research should consider testing these devices to better reflect real-world conditions. We also suggest exploring other physiological signals, such as eye gaze or facial expressions. Combining these with cognitive effort indicators, like changes in $\Delta HbO$ and user experience scores, could offer deeper insights. Our study focused only on hand and stylus input. While these are common in touchscreen learning tools, we did not include other inputs like a mouse, keyboard, or virtual reality. These are widely used and relevant in both traditional and immersive learning environments. Their exclusion limits how broadly our results can be applied.

### v6

@@ Abstract.
Interaction system; touch-based interaction; stylus; functional Near-Infrared Spectroscopy (fNIRS); brain signal; hemodynamic response; touchscreens; educational games
journalyear: 2024copyright: rightsretainedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: 27th International Conference on Multimodal Interaction (ICMI 2025) ; October 13–17, 2025; Canberra, Australiaisbn: 978-1-4503-XXXX-X/2018/06ccs: Human-centered computing Human computer interaction (HCI)ccs: Applied computing Interactive learning environmentsccs: Applied computing Learning management systemsccs: Applied computing Computer-assisted instruction
Figure 1. The overview of the study: (a) Experimental setup: A participant plays an educational quiz game on a touchscreen laptop while functional near-infrared spectroscopy (fNIRS) records prefrontal hemodynamic activity. (b) Educational Game Environment: Sample game interface with multiple-choice questions. (c) Hemodynamic Response: Heatmap shows significantly lower oxygen demand when using the hand vs. stylus (d) Cognitive effort: Hand requires significantly less cognitive effort despite similar performance , which indicates neural efficiency differences by input modality.
@@ 1. Introduction
Touch-based interaction is becoming a more powerful modality due to its resemblance to the traditional pen-and-paper experience.  (Elttayef, 2016; Li et al., 2013; Taranta et al., 2016). Incorporating touchscreens into educational games has changed the way we interact while playing. This offers easy touch and movement on the screen, and makes applications more user friendly  (Crescenzi-Lanna, 2020). Traditional human-computer interaction evaluations, such as questionnaires and think-aloud protocols or user behavior metrics such as reaction times, have limitations, including susceptibility to ambiguity, social pressures, and disruption of the interaction  (Frey et al., 2016; Kosch et al., 2023; Pettersson et al., 2018; Yu et al., 2023). The real-time behavioral metrics such as an functional near-infrared spectroscopy (fNIRS) data offer quantitative insights  (Lloyd-Fox et al., 2010; Yücel et al., 2021) where the traditional methods fail to accurately represent users’ mental states as they are unable to discern between factors such as low concentration or task complexity  (Frey et al., 2016; Fairclough, 2009; Frey et al., 2013; Pike et al., 2012) and attention as well as motivation  (Pike et al., 2012).
fNIRS, a noninvasive optical imaging technology, measures the changes in concentrations of oxygenated hemoglobin ($\Delta HbO$) and deoxygenated hemoglobin in the prefrontal cortex (PFC) to evaluate hemodynamic responses  (Lloyd-Fox et al., 2010; Herff et al., 2014; Yücel et al., 2021; Shewokis et al., 2015; Getchell and Shewokis, 2023; Yu et al., 2023; Sharmin and Barmaki, 2024; Sharmin et al., 2024b). The fNIRS information provides human computer interaction researchers with a powerful tool to explore the complex relationship between cognitive processes and interactive technologies  (Frey et al., 2016; Kosch et al., 2023; Pettersson et al., 2018; Sharmin and Barmaki, 2024). This method assumes that brain or neural activity leads to an increase in oxygenated blood flow, as indicated by increased $\Delta HbO$. This suggests that during task performance, it requires more cognitive effort  (Sharmin et al., 2024b; Lloyd-Fox et al., 2010; Herff et al., 2014; Shewokis et al., 2015; Yücel et al., 2021; Kosch et al., 2023; Pettersson et al., 2018; Yu et al., 2023). This connection helps us to understand the brain mechanisms behind cognitive functions, and build technologies and therapies to optimize or monitor them  (Lloyd-Fox et al., 2010; Yücel et al., 2021). The oxygen flow to the PFC can also be used to measure relative neural efficiency, which quantifies cognitive effort and behavioral performance. It can also measure relative neural involvement, which illustrates the relationship among motivation, engagement, and performance  (Paas et al., 2003; Shewokis et al., 2015; Getchell and Shewokis, 2023).
@@ 3.1. Participants
Outliers were identified and removed based on the standard interquartile range (IQR) method ($Q1-1.5\times IQR$ or $Q3+1.5\times IQR$)(Dodge, 2008). If the average $\Delta HbO$ values of a participant fell outside the range, they were considered outliers. The final sample comprised 16 graduate students (11 female, 5 male), aged 23–32 years ($M=27.13\pm 2.5$). Thus, we removed two more outliers after data collection. Participants were graduate students aged between 23 and 32 years (11F,$M=27.13\pm 2.5$). The participants had strong experience using their hands on touchscreens, but only a few were confident using a stylus ($18.75\%$). More than half had never used games as a learning tool, while others used them occasionally. In terms of graph understanding, only a small number (6.25%) felt somewhat confident.
@@ 3.2. Apparatus
In this study, we used two computers and an fNIRS device (Figure 1 (a)). We used an 18-channel fNIRS imager 2000S (fNIRS Device LLC, Potomac, MD, USA), four light emitters with wavelengths ($\lambda$) of 730-850 nm and an average power of $<1$ mW, sampled at 2 Hz. The detectors were separated by 2.5 cm, which resulted in a penetration depth of approximately 1.2 cm. The fNIRS sensor band was aligned to match the horizontal and vertical axes of the head with those of the band. Specifically, the sensor’s vertical axis was placed at the Fp1 and Fp2 locations delineated in the international 10 - 20 system of cerebral electrode placement. Data was converted to concentration changes using the modified Beer-Lambert law (mBLL). The educational quiz games were played on a Dell XPS touchscreen laptop (L1) [Intel (R) Core(TM) i7-7700HQ processor at 2.80 GHz]. This laptop supports both hand touch and stylus input with ten touch points. We used a desktop computer (L2) [Intel(R) Core(TM) i7-10700T processor at 2.00 GHz] connected to the fNIRS device. Cognitive optical brain imaging studio Software was installed on L2 to gather fNIRS data. fNIRSoft Software (Version 4.9) evaluated the data. A Python code was executed on L2 and synchronized with each task to signal the fNIRS device to differentiate between the quiz and rest periods (Sharmin and Abrar, 2025; Sharmin et al., 2024a).
@@ 4.1. Quantitative Analysis
    These values were converted to Z-scores to standardize the data. Our aim was better performance with less cognitive effort. We used the mean task performance score to calculate the standardized performance index ($P_{z}$) as shown in Equation 2. We also computed the inverse of the mean $\Delta HbO$ values to reflect reduced effort, and used this in calculating the standardized average oxygenated hemoglobin index ($M_{z}$), as shown in Equation 3 (Aksoy et al., 2025; Paas et al., 2003; Shewokis et al., 2015; Getchell and Shewokis, 2023; Koiler et al., 2022). In the X-axis, we plot the standardized average $\Delta HbO$ and on the Y-axis, we plot the standardized performance level $P_{z}$. Then, we used a Euclidean transformation from vector geometry. The equation for RNE in Equation 4 and RNI in Equation 5 represent projections of the vector onto two rotated coordinate axes: one axis captures efficiency (RNE line, running bottom-left to top-right) and the other captures involvement (RNI line, bottom-right to top-left). These rotated axes help to separate two cognitive dimensions: how effectively someone learns, and how deeply involved they are.
@@ 4.2. Self-reported Measures & Instruments
We used the System Usability Scale (SUS) (Brooke et al., 1996), ten questions with a 5-point Likert-scale ranging from ‘strongly disagree’ to ‘strongly agree.’ On a scale of 0 to 100, the final SUS score was determined (0–50 = not acceptable, 51–67 = poor, 68 = okay, 69–80 = good, 81–100 = exceptional) (Brooke et al., 1996). We also used the NASA-TLX, which consists of six subscales: mental demand, physical demand, temporal demand, performance, effort, and frustration, to assess the mental workload of participants during learning procedures. This assessment has six questions to produce an overall workload rating of 0-100 (Hart and Staveland, 1988).
@@ 5.1.2. Hemodynamic Response
Figure 5. Heatmap of $\Delta HbO$ across four regions of the PFC: (a) hand condition and (b) stylus condition. The stylus condition exhibited overall significantly higher $\Delta HbO$ which signifies greater neural activity compared to the hand condition ($p<0.05$) across the PFC.
@@ 5.1.3. Relative Neural Efficiency (RNE) and Relative Neural Involvement (RNI)
Figure 6. Relative neural efficiency and involvement: (a) relative neural efficiency (RNE), where the hand condition shows significantly higher neural efficiency ($p<0.05$) with lower cognitive effort, and (b) lower Relative Neural Involvement (RNI) compared to the stylus condition ($p<0.05$). The red dot represents the hand, and the green diamond represents the stylus condition. The cognitive effort is higher in the stylus although the performance level is almost the same.
@@ 6. Discussion
We investigated how input modalities (hand vs. stylus) impact neural activity, cognitive effort, and user experience in an educational game. Our findings show no significant difference in performance between modalities, suggesting both are equally effective for simple tap-based quiz tasks.
@@
$RQ_{1}$ focused on the hemodynamic response in educational gaming with two input methods (hand and stylus). Higher $\Delta HbO$ suggests higher oxygen flow to the brain, which is the result of higher neural activity (Kerr et al., 2022; Sharmin et al., 2024b). In our study, the change in blood oxygenation levels ($\Delta HbO$) showed a significant decrement with hand usage compared to the stylus. The higher $\Delta HbO$ while using a stylus suggests that handling a stylus can require more precise motor control, fine motor skills, and coordination than using the hand directly. Thus, it is observed that the type of input method may change the cognitive effort in educational game studies. According to the demographic data, most participants were not regular stylus users. While this unfamiliarity could have potentially influenced their performance by increasing cognitive effort, our data did not reflect any significant disruption. Both response time and performance scores did not show any significant differences. This indicates that prior exposure to the stylus did not negatively affect their task performance, particularly for simple actions like tapping. However, this also highlights a critical design consideration: learning materials and interactive tools should be designed in such a way that users invest their effort primarily in the learning process, not in struggling to operate the interface itself.
@@
$RQ_{2}$ focused on observing cognitive effort by measuring relative neural efficiency (RNE) and relative neural involvement (RNI). The results suggest that using the hand required less cognitive effort, as reflected in higher RNE and lower RNI values. On the other hand, the stylus condition showed higher neural involvement, but the performance score was almost the same as the hand condition. In the stylus condition, participants had to manage both holding the stylus and responding to quiz questions simultaneously, which required more tasks and led to more cognitive effort (Schwarz et al., 2014; Koiler et al., 2022; Getchell and Shewokis, 2023). This suggests that hands are more natural or intuitive in educational game play and also highlights the importance of interaction methods that support efficient performance with lower cognitive effort.
@@
$RQ_{3}$ focused on the impact of self-reported usability and task load measures of different input methods in playing educational games. The SUS score showed that users found the hand-based input more intuitive and usable, with a good usability score. The NASA-TLX results indicated that the hand condition required less effort and caused less frustration. Together, these findings suggest that hand-based input is more efficient and less demanding cognitively than stylus-based interaction for educational gameplay.
@@
According to the results of $RQ_{1}$, $RQ_{2}$, and $RQ_{3}$, the hand condition provides better outcomes than the stylus in both objective and subjective evaluations. Lower $\Delta HbO$ levels with hand usage are revealed by hemodynamic data ($RQ_{1}$), which suggests less cognitive load. This is further supported by cognitive effort measurements ($RQ_{2}$). These results suggest that the natural and intuitive usage of the hand leads to higher efficiency (RNE) and lower involvement (RNI) than stylus usage. These findings are supported by self-reported measures ($RQ_{3}$), as NASA-TLX results show less task load and annoyance, and SUS ratings show improved usability for hand input. According to these findings, it might be easier to use the hand as an input method in educational games. The degree to which subjective impressions and objective facts agree demonstrates the importance of taking both viewpoints into account when assessing interaction techniques. This highlights the possibilities of hand-based input in learning games and provides information for developing systems that try to reduce cognitive effort while increasing effectiveness and user satisfaction.
@@
Our findings have broader implications for educational technology and natural user interface design. The lower cognitive effort in the hand condition shows that familiarity and intuitiveness are important for user experience and performance. These results contribute to human-computer interaction by showing that neuronal efficiency improves when input complexity is reduced. The study also shows that neurophysiological data, like fNIRS, can validate design choices that are often based on subjective input. By showing cognitive load differences between hand and stylus, our work supports a user-centered approach to input selection. This neuro-informed view can guide future studies on input methods like mouse and interface designs, and help develop adaptive educational tools that match interaction styles to users’ cognitive and behavioral performance.
Limitations and Future Work:
@@
This study may have been limited by its small sample size. Differences in participants’ age, education, and background may have influenced the results. Future studies will include a more diverse and balanced group. A laptop setup may have reduced natural interaction; using a tablet or iPad may create a more intuitive experience. Future work should test these devices and explore real-world educational tasks such as drawing, dragging, or multitasking. Other physiological signals, such as eye gaze or facial expressions, should also be explored. Combining these with $\Delta HbO$ and user experience scores may provide deeper insights. This study focused only on hand and stylus input. We did not include mouse, keyboard, or virtual reality, which are common in learning environments. Their exclusion limits the generalizability of our results.